She needs to prepare 24 ounces of a 10​% hydrochloric acid solution. Find the amount of 12​% solution and the amount of 6​% solution she should mix to get this solution
Answer by fractalier(6550)

Let's call the amount of 12% used, x.
Then the amount of 6% used would have to be 24-x (since we need 24 ounces total).
Here's the set up...
.12x + .06(24-x) = .10(24)
Now solve for x...
.12x + 1.44 - .06x = 2.4
.06x = .96
x = 16 ounces of 12%
24-x = 8 ounces of 6%